### v1

@@ 2.3.1 Rotation
One method of inducing rotation in the system is to stir the quantum fluid, typically carried out by a moving barrier [44, 46, 79, 47]. It should be emphasized however that such an operation lead to rotations following a quite different mechanism to stirring a spoon in a classical liquid (that in fact would not set a superfluid in motion). Indeed, the quantum gas is stirred by producing a dark soliton-type, namely a spatial region in the annulus in which the condensed fraction get suppressed to a point where phase slips can occur with negligible energy cost. By moving the dark soliton-type, to compensate for the phase kinking or slippage in the depletion region, the condensate wavefunction develops a nonzero phase gradient ultimately putting the fluid in motion. Such a protocol can be realized, for example, by driving the condensate with a focused blue-detuned laser beam. In what follows, we demonstrate how to instigate a persistent current flow in a cold atoms platform through stirring.
@@
In order to rotate the condensate, we place a time-dependent potential barrier $V(x-\Omega Rt)$ that moves at an angular velocity $\Omega$, such that the total system is described by this Hamiltonian $\mathcal{H}(\Omega,t)=\mathcal{H}_{0}+V(x-\Omega Rt)$. The barrier is assumed to be switched on adiabatically, so no high-energy excitations are generated in the condensate. By performing a suitable unitary transformation $\mathcal{H}_{rot}=\mathcal{U}^{\dagger}(t)\mathcal{H}(\Omega,t)\mathcal{U}(t)$ with $\mathcal{U}=\mathrm{exp}(\imath L_{z}\Omega t/\hbar)$, we can switch over to the rotating reference frame with a frequency $\Omega$ as that of the potential barrier, thereby eliminating the time dependence of the Hamiltonian leaving us with $\mathcal{H}_{rot}=\mathcal{H}_{0}+V(x)-\Omega\hat{L}_{z}$, where $L_{z}$ is the $z$-component of the angular momentum perpendicular to the $x$-$y$ plane in which the ring is taken to lie in [44, 173]. Translating to polar coordinates as in Sec. 2.1 and expressing the angular momentum as $-\Omega L_{z}=\imath\hbar\Omega\frac{\partial}{\partial\varphi}$, the Hamiltonian can be written as
@@
The presence of the rotating barrier introduces a shift in the momentum operator and, in turn, in the system’s total energy. Suppose we ignore the interaction and look at the expression for the single-particle case, it is clear that the Hamiltonians of Eqs. (1) and (26) are of a similar qualitative structure by choosing a gauge field $\chi\mathbf{A}=m\Omega R^{2}$  [^7] Note that we take the gauge in (1) to be zero and ignore the last term in (26) as this corresponds to a uniform shift in the energy spectrum.. Therefore, the behaviour of a charged particle in a magnetic field can be emulated in a cold atoms setting by incorporating a synthetic gauge field $m\Omega R$. The analogy can be extended even further to the exerted forces, with the Lorentz force being mimicked by the Coriolis force, which originates as a response to the induced rotation.
@@
which at a specific site $j$ is the difference between the incoming and outgoing particles. In contrast to the rotating frame where the current is a sawtooth, the one in the lab frame is given by discrete steps – see Fig. 3(b). The ideal Bose gas is a special case. Due to the spectrum of the energy against the total momenta, the ground-state of the system being at $\ell=0$, the current obtained is not metastable as it decays one the source imparting rotation is removed. For this reason, we showcase the spectrum for weakly repulsive interactions ($c=10$) by obtaining the momenta through Bethe ansatz, to be discussed in more detail in Sec. 3.1.5. The plateaus reflect the quantized nature of the angular momentum associated with the circulating particles, which increases on each subsequent jump [^8] In the case of fermions, the parity of the current is captured through these jumps, occurring at integer of half-odd integer values of the flux.. It must be stressed that in the non-rotating frame, we get a better picture of the trade-off between the stability of the system and the magnitude of the current (circulation values) through the energy of the system (see Fig. 3(a)). To offset the instability of the system at higher rotation rates, one expects the formation of excitations, such as vortices, that reduce the total energy [46, 174].
@@ 2.3.4 Machine learning
For atomic currents, the first theoretical and numerical study has been done in Ref. [200]. The authors propose to improve the generation of persistent currents using deep reinforcement learning. The learning setup is shown in Fig. 4.
Figure 4: Deep reinforcement learning to optimize the preparation protocol for generating atomic current states in ring lattices. (a) Deep neural network which suggests lattice potentials for each timestep of the protocol. Neural network is trained using observables from the quantum system. (b) Atomic ring lattice where suggested potential is applied to create current state. The final state is measured and observables such as the momentum distribution are used to train neural network (c) Optimised protocol for potential to generate current, showing lattice potential for different lattice sites $m$ and discrete timesteps $t_{n}$. Figure from Ref. [200].
A goal is to find a fast protocol that create a specific atomic current state in an atomic ring lattice. In such a ring lattice, one can control the potential of each lattice site in time. Experimentally, this can be achieved by time-dependent control of the potential of the ring lattice by light-shaping techniques [201]. By changing the potential in a controlled way, specific current states can be generated. The authors find specific protocols that generate such currents in very short times using reinforcement learning [193]. Here, a machine learning algorithm directly controls the protocol to generate currents and learns from experience to achieve the target current. The particular method used is called proximal policy learning [202] where two deep neural network, called actor and critic, work together to find good protocols. The actor network decides on the potential to choose at each timestep, while the critic evaluates the former’s decisions by assigning a score. After the protocol chosen by the actor has been run, the current of the system is measured and a reward is given to the neural network depending on how close the actual value is to the desired one. The weights of the network are adjusted depending on the reward. This learning process is repeated until it converges to a good solution.
@@ 2.3.5 Experimental methods
Alternatively, it is possible to use two optical fields to couple between two different internal states in what is referred to as a Raman process. In this case, it is possible for the levels to be energetically resolved without a large momentum transfer, making it possible to use collinearly propagating beams. As long as the energy difference between the two selected levels is not too large (as is generally true for coupling between different magnetic sublevels levels in the electronic ground state) the linear momentum imparted to the atoms by a single collinear pulse pair is small enough to be neglected. This Raman-beam approach to orbital angular momentum control of a BEC was first demonstrated by Wright et al. [206] with a simply-connected BEC, and was later used for some of the persistent current experiments with ring-BECs at NIST [49] and at Cambridge [50].
There are some important factors to be considered in the use of these kinds of two-photon techniques. With suitable laser intensities and moderate detunings from the $D$ lines of alkali metal atoms, the $\pi$-pulse times can be made as short as a few microseconds, but attention must be paid to keeping the single-photon scattering rate low enough to prevent unacceptable levels of heating and loss. A second challenge in using two-photon techniques is that it not generally safe to ignore the detuning and state-dependent light shifts that occur when applying such fields, especially for single-photon detunings on the order of the hyperfine splittings. Finally, any realistic experimental design must account for the possibility of spatially-varying Rabi frequencies and light shifts, because the phase singularities present in laser fields with nonzero orbital angular momentum necessarily make them spatially inhomogeneous [206]. These drawbacks were less important in the earliest experiments [49, 50], when techniques for detecting the circulation state of the ring were less well developed, and it was advantageous to use a technique where the atomic orbital angular momentum was guaranteed to change by a quantized amount, as long as it reached the target state.
@@
Techniques for creating persistent currents with a single far-detuned single-frequency laser field are comparatively simpler in many ways, and they have become more commonly used in recent experiments. Here, we will separately consider the two very different approaches of stirring and phase imprinting. The idea of stirring a superfluid to induce circulation is conceptually simple: create a localized disturbance in the potential and create a flow by moving it faster than the superfluid critical velocity. We note that quantum phase slips dynamics across the potential disturbance determines the final winding number of the superfluid flow –see Sec. 4. In practice, there are many details about the creation and evolution of various types of excitations (phonons, solitons, vortices) in the superfluid that can affect the success and reliability of such an approach. One has the choice of whether to apply a repulsive or attractive potential defect, and its size and shape compared to the channel width. There are also questions of how to turn it on, how it should move, and how to turn it off. Which technique is optimal for a given situation is highly dependent on system characteristics and experimental goals.
The first experiment creating a persistent current in a ring BEC via a stirring technique used a repulsive (blue detuned optical) barrier that was thin in the azimuthal direction and large enough radially to span the width of the ring, as shown in Fig. 6 taken from Ref. [46]. The position of the spot was controlled using a 2D acousto-optic deflector, which allowed for the use of shaped, time-averaged potentials by rapidly scanning a tightly focused spot over a defined trajectory. In the case of some of the early ring BEC work done at LANL, the ring trap was a time-averaged “painted” potential generated by a scanned red-detuned beam, and movable potential barriers were formed by shutting off the scanned beam at one or more points, instead of by adding a repulsive beam [132].
@@
The first experimental implementation of this technique was in the work done at LENS by Del Pace et al [68]. Because that work was done using a superfluid of fermionic 6Li atoms they were able to investigate how the response to phase imprinting changed with interaction strength, with some interesting differences observed between the BEC, unitary, and BCS regimes. Most importantly, they found that it was possible to reliably imprint a phase winding on the superfluid ring, even without modifying the density profile with a repulsive barrier as had been suggested in previous theoretical work. The dynamics at the phase dislocation were found to be important, leading to differences in the (higher chemical potential) unitary and BCS regimes, where it was necessary to apply a larger phase gradient to achieve a given winding number in the superfluid (See Fig. 7). Inserting a repulsive optical barrier at the location of the gradient discontinuity reduced the unwanted density excitations as expected, but was found to degrade the reliability of initializing the system in the $w=1$ state. With up to 5 repeated applications of the phase imprinting pulse, they were also able to realize higher winding number states of up to $w=9$. Further details of this work pertaining to lifetime and decay will be described in Sec. 5.2.
Before concluding this section, we point out that essentially any method that has been used to generate vortices in a simply-connected system can potentially be adapted to generate a persistent current around an annulus. The technique of magnetic phase imprinting was proposed by Nakahara et al. [210] and was first implemented by the Ketterle group [211] to generate vortices. It involves adiabatically inverting the direction of the axial part of a magnetic (or hybrid magnetic-optical) type trap, which creates a cross disgyration in the BEC as the axial field goes through zero. If the atoms are polarized in a single internal spin state with projection $m$ on the local field, and they adiabatically follow the field reversal during the transformation, the winding number of the current must necessarily change by $2m$. While this technique has not yet been used with a toroidally trapped condensate, it remains a possibility for the future.
@@ 3 Theoretical framework
This section covers the main theoretical tools used in the forthcoming parts of the review. We discuss the one-dimensional limit of bosonic and fermionic quantum fluids, which corresponds to simple models where persistent currents have been studied and where exact solutions are available. Since the flow dynamics along the ring is quasi-1D, we have chosen to focus mainly on the one-dimensional treatments, which account for several important features of the system as flux quantization, phase slips, and structure of the read-out interference patterns. We also cover experimentally relevant situations of annular traps having several radial modes occupied, which are described by a three-dimensional model at weak interactions. In particular, this model allows us to cover the emergence of structures in the transverse size of the ring, such as e.g. vortices. Examples of both 1D and 3D dynamics will be provided in the forthcoming Sec. 4.
@@ 3.1 Bosons
Bosonic particles are characterized by having a wavefunction completely symmetric under particle exchange. This important property ensures the possibility of occupying the same single-particle state in the case of non-interacting bosons, leading to Bose-Einstein condensation, which corresponds to the creation of a giant matter-wave. Condensates display off-diagonal long-range order (ODLRO), that corresponds to large-distance phase coherence. In the case of interacting bosonic fluids, quantum depletion of the condensate state usually occurs. Still, they display the phenomomenon of superfluidity, which may occur even in the absence of Bose-Einstein condensation. In low dimensions, phase fluctuations and correlations are enhanced, and coherence is degraded especially at large interactions. In this case, only quasi-ODLRO is possible. In this section, we present the main theoretical tools that we have at hand to describe interacting Bose fluids in a ring geometry, ranging from mean-field approaches to effective field theories and lattice models.
@@ 3.1.1 Mean-field theories
Starting from Eq. (3.1.1) one can derive a simpler description of the bosonic particles in the limit of large $N_{p}$, weak interactions and zero temperatures, where, to a very good approximation, all bosons macroscopically occupy the same quantum state. At these low temperatures and for a large enough sample, the system becomes very diluted and one can simplify the generic interactions to be only two-body $s$-wave collisions. The interacting potential can then be approximated by a pseudo-potential description $\mathcal{V}(\mathbf{r}-\mathbf{r^{\prime}})=g_{3D}\delta(\mathbf{r}-\mathbf{r^{\prime}})$ with strength $g_{3D}=4\pi^{2}a_{s}/m$ characterized by the $s$-wave scattering length, $a_{s}$, which can be positive ($a_{s}>0$) or negative ($a_{s}<0$) indicating that the effective interactions are repulsive or attractive, respectively. These scattering lengths can, depending on the atomic species, be tuned in magnitude and sign using magnetic [212, 213] or optical [214, 215] Feshbach resonances.

### v3

@@ 2.3.1 Rotation
One method of inducing rotation in the system is to stir the quantum fluid, typically carried out by a moving barrier [48, 50, 82, 51]. It should be emphasized however that such an operation leads to rotations following a quite different mechanism to stirring a spoon in a classical liquid (that in fact would not set a superfluid in motion). Indeed, the quantum gas is stirred by producing a dark soliton-type, namely a spatial region in the annulus in which the condensed fraction gets suppressed to a point where phase slips can occur with negligible energy cost. By moving the dark soliton-type, to compensate for the phase kinking or slippage in the depletion region, the condensate wavefunction develops a nonzero phase gradient ultimately putting the fluid in motion. Such a protocol can be realized, for example, by driving the condensate with a focused blue-detuned laser beam (see Fig. 6 for an example). In what follows, we demonstrate how to instigate a persistent current flow in a cold atoms platform through stirring.
@@
In order to rotate the condensate, we place a time-dependent potential barrier $V(x-\Omega Rt)$ that moves at an angular velocity $\Omega$, such that the total system is described by this Hamiltonian $\mathcal{H}(\Omega,t)=\mathcal{H}_{0}+V(x-\Omega Rt)$. The barrier is assumed to be switched on adiabatically, so no high-energy excitations are generated in the condensate. By performing a suitable unitary transformation $\mathcal{H}_{rot}=\mathcal{U}^{\dagger}(t)\mathcal{H}(\Omega,t)\mathcal{U}(t)+\imath\partial_{t}\mathcal{U}(t)$ with $\mathcal{U}=\mathrm{exp}(\imath L_{z}\Omega t/\hbar)$, we can switch over to the rotating reference frame with a frequency $\Omega$ as that of the potential barrier, thereby eliminating the time dependence of the Hamiltonian leaving us with $\mathcal{H}_{rot}=\mathcal{H}_{0}+V(x)-\Omega\hat{L}_{z}$, where $L_{z}$ is the $z$-component of the angular momentum perpendicular to the $x$-$y$ plane in which the ring is taken to lie in [48, 193]. Translating to polar coordinates as in Sec. 2.1 and expressing the angular momentum as $-\Omega L_{z}=\imath\hbar\Omega\frac{\partial}{\partial\varphi}$, the Hamiltonian can be written as
@@
The presence of the rotating barrier introduces a shift in the momentum operator and, in turn, in the system’s total energy. Suppose we ignore the interaction and look at the expression for the single-particle case, it is clear that the Hamiltonians of Eqs. (1) and (23) are of a similar qualitative structure by choosing a gauge field(4)(4)(4)Note that we take the gauge in (1) to be zero and ignore the last term in (23) as this corresponds to a uniform shift in the energy spectrum. $\chi\mathbf{A}=m\Omega R^{2}$.Therefore, the behaviour of a charged particle in a magnetic field can be emulated in a cold atoms setting by incorporating a synthetic gauge field $m\Omega R$. The analogy can be extended even further to the exerted forces, with the Lorentz force being mimicked by the Coriolis force, which originates as a response to the induced rotation.
@@
which at a specific site $j$ is the difference between the incoming and outgoing particles. In contrast to the rotating frame where the current is a sawtooth, the one in the lab frame is given by discrete steps – see Fig. 3(b). The ideal Bose gas is a special case. Due to the spectrum of the energy against the total momenta, the ground-state of the system being at $\ell=0$, the current obtained is not metastable as it decays once the source imparting rotation is removed. For this reason, we showcase the spectrum for weakly repulsive interactions ($c=10$) by obtaining the momenta through Bethe ansatz, to be discussed in more detail in Sec. 3.1.5. The plateaus reflect the quantized nature of the angular momentum associated with the circulating particles, which increases on each subsequent jump.(5)(5)(5)In the case of fermions, the parity of the current is captured through these jumps, occurring at integer of half-odd integer values of the flux. It must be stressed that in the non-rotating frame, we get a better picture of the trade-off between the stability of the system and the magnitude of the current (circulation values) through the energy of the system (see Fig. 3(a)). To offset the instability of the system at higher rotation rates, one expects the formation of excitations, such as vortices, that reduce the total energy [50, 194].
@@ 2.3.4 Machine learning
For atomic currents, the first theoretical and numerical study has been done in Ref. [220]. The authors propose to improve the generation of persistent currents using deep reinforcement learning. An example of this learning setup is shown in Fig. 4.
Figure 4: An example for deep reinforcement learning to optimize the preparation protocol for generating atomic current states in ring lattices. (a) Deep neural network which suggests lattice potentials for each timestep of the protocol. Neural network is trained using observables from the quantum system. (b) Atomic ring lattice where suggested potential is applied to create current state. The final state is measured and observables such as the momentum distribution are used to train neural network. (c) Optimised protocol for potential to generate current, showing lattice potential for different lattice sites $m$ and discrete timesteps $t_{n}$. Figure from Ref. [220].
A goal is to find a fast protocol that creates a specific atomic current state in an atomic ring lattice. In such a ring lattice, one can control the potential of each lattice site in time. Experimentally, this can be achieved by time-dependent control of the potential of the ring lattice by light-shaping techniques [221]. By changing the potential in a controlled way, specific current states can be generated. The authors find specific protocols that generate such currents in very short times using reinforcement learning [213]. Here, a machine learning algorithm directly controls the protocol to generate currents and learns from experience to achieve the target current. The particular method used is called proximal policy learning [222] where two deep neural networks, called actor and critic, work together to find good protocols. The actor network decides on the potential to choose at each timestep, while the critic evaluates the former’s decisions by assigning a score. After the protocol chosen by the actor has been run, the current of the system is measured and a reward is given to the neural network depending on how close the actual value is to the desired one. The weights of the network are adjusted depending on the reward. This learning process is repeated until it converges to a good solution.
@@ 2.3.5 Experimental methods
Alternatively, it is possible to use two optical fields to couple between two different internal states in what is referred to as a Raman process. In this case, it is possible for the levels to be energetically resolved without a large momentum transfer, making it possible to use collinearly propagating beams. As long as the energy difference between the two selected levels is not too large (as is generally true for coupling between different magnetic sublevels in the electronic ground state) the linear momentum imparted to the atoms by a single collinear pulse pair is small enough to be neglected. This Raman-beam approach to orbital angular momentum control of a BEC was first demonstrated by Wright et al. [226] with a simply-connected BEC, and was later used for some of the persistent current experiments with ring-BECs at NIST [53] and at Cambridge [54].
There are some important factors to be considered in the use of these kinds of two-photon techniques. With suitable laser intensities and moderate detunings from the $D$ lines of alkali metal atoms, the $\pi$-pulse times can be made as short as a few microseconds, but attention must be paid to keeping the single-photon scattering rate low enough to prevent unacceptable levels of heating and loss. A second challenge in using two-photon techniques is that it not generally safe to ignore the detuning and state-dependent AC Stark shifts that occur when applying such fields, especially for single-photon detunings on the order of the hyperfine splittings. Finally, any realistic experimental design must account for the possibility of spatially-varying Rabi frequencies and AC Stark shifts, because the phase singularities present in laser fields with nonzero orbital angular momentum necessarily make them spatially inhomogeneous [226]. These drawbacks were less important in the earliest experiments [53, 54], when techniques for detecting the circulation state of the ring were less well developed, and it was advantageous to use a technique where the atomic orbital angular momentum was guaranteed to change by a quantized amount, as long as it reached the target state.
@@
Techniques for creating persistent currents with a single far-detuned single-frequency laser field are comparatively simpler in many ways, and they have become more commonly used in recent experiments. Here, we will separately consider the two very different approaches of stirring and phase imprinting. The idea of stirring a superfluid to induce circulation is conceptually simple: create a localized disturbance in the potential and create a flow by moving it faster than the superfluid critical velocity. We note that quantum phase slips dynamics across the potential disturbance determines the final winding number of the superfluid flow - see Sec. 4. In practice, there are many details about the creation and evolution of various types of excitations (phonons, solitons, vortices) in the superfluid that can affect the success and reliability of such an approach. One has the choice of whether to apply a repulsive or attractive potential defect, and its size and shape compared to the channel width. There are also questions of how to turn it on, how it should move, and how to turn it off. Which technique is optimal for a given situation is highly dependent on system characteristics and experimental goals.
The first experiment creating a persistent current in a ring BEC via a stirring technique used a repulsive (blue detuned optical) barrier that was thin in the azimuthal direction and large enough radially to span the width of the ring, as shown in Fig. 6 taken from Ref. [50]. The position of the spot was controlled using a 2D acousto-optic deflector, which allowed for the use of shaped, time-averaged potentials by rapidly scanning a tightly focused spot over a defined trajectory. In the case of some of the early ring BEC work done at Los Alamos National laboratory (LANL), the ring trap was a time-averaged “painted” potential generated by a scanned red-detuned beam, and movable potential barriers were formed by shutting off the scanned beam at one or more points, instead of by adding a repulsive beam [138].
@@
The first experimental implementation of this technique was in the work done at LENS by del Pace et al [57]. Because that work was done using a superfluid of fermionic 6Li atoms they were able to investigate how the response to phase imprinting changed with interaction strength, with some interesting differences observed between the BEC, unitary, and BCS regimes. Most importantly, they found that it was possible to reliably imprint a phase winding on the superfluid ring, even without modifying the density profile with a repulsive barrier as had been suggested in previous theoretical work. The dynamics at the phase dislocation were found to be important, leading to differences in the (higher chemical potential) unitary and BCS regimes, where it was necessary to apply a larger phase gradient to achieve a given winding number in the superfluid (see Fig. 7). Inserting a repulsive optical barrier at the location of the gradient discontinuity reduced the unwanted density excitations as expected, but was found to degrade the reliability of initializing the system in the $w=1$ state. With up to 5 repeated applications of the phase imprinting pulse, they were also able to realize higher winding number states of up to $w=9$. Further details of this work pertaining to lifetime and decay will be described in section 5.2.
Before concluding this section, we point out that essentially any method that has been used to generate vortices in a simply-connected system can potentially be adapted to generate a persistent current around an annulus. The technique of magnetic phase imprinting was proposed by Isoshima et al. [230] and was first implemented by the Ketterle group [231] to generate vortices. It involves adiabatically inverting the direction of the axial part of a magnetic (or hybrid magnetic-optical) type trap, which creates a cross disgyration in the BEC as the axial field goes through zero. If the atoms are polarized in a single internal spin state with projection $m$ on the local field, and they adiabatically follow the field reversal during the transformation, the winding number of the current must necessarily change by $2m$. While this technique has not yet been used with a toroidally trapped condensate, it remains a possibility for the future.
@@ 3 Theoretical framework
This section covers the main theoretical tools needed to understand and characterize persistent currents in both bosonic and fermionic gases. We start from the mean-field model for bosons, which can describe weakly interacting condensates confined in annular traps with several radial and axial modes occupied. We then focus on the one-dimensional limit corresponding to very tight transverse trapping confinement both in radial and axial direction. While early experimental results can be described by mean-field, quantum fluctuations are expected to become increasingly relevant in future experiments. In this regime, we introduce the theoretical frameworks allowing us to account for these quantum fluctuations and beyond mean-field effects, both in an exact way as well as through effective quantum field theory. We then turn to fermionic systems and present the main formalism needed to describe the BCS-BEC crossover in Fermi gases. Also for fermionic systems, we then focus on the one-dimensional limit, in particular on the SU(2) and SU(N) Fermi Hubbard models, where striking beyond-mean field effects emerge in the persistent currents. We finally present and dwell on the theory for the readout of the persistent currents in cold atomic systems through interference patterns. Examples of both 1D and 3D dynamics will be provided in the forthcoming Secs. 4 and 5.
@@ 3.1 Bosons
Bosonic particles are characterized by having a wavefunction completely symmetric under particle exchange. This important property ensures the possibility of occupying the same single-particle state in the case of non-interacting bosons, leading to Bose-Einstein condensation, which corresponds to the creation of a giant matter-wave. Condensates display off-diagonal long-range order (ODLRO), that corresponds to large-distance phase coherence. In the case of interacting bosonic fluids, quantum depletion of the condensate state usually occurs. Still, they display the phenomenon of superfluidity, which may occur even in the absence of Bose-Einstein condensation -see Sec. 2.2. In low dimensions, phase fluctuations and correlations are enhanced, and coherence is degraded especially at large interactions. In this case, only quasi-ODLRO is possible. In this section, we present the main theoretical tools that we have at hand to describe interacting Bose fluids in a ring geometry, ranging from mean-field approaches to effective field theories and lattice models.
@@ 3.1.1 Mean-field theories
Starting from Eq. (3.1.1) one can derive a simpler description of the bosonic particles in the limit of large particle number $N_{p}$, weak interactions and zero temperatures, where, to a very good approximation, all bosons macroscopically occupy the same quantum state. At these low temperatures and for a large enough sample, the system becomes very diluted and one can simplify the generic interactions to be only two-body $s$-wave collisions. The interacting potential can then be approximated by a pseudo-potential description $\mathcal{V}(\mathbf{r}-\mathbf{r^{\prime}})=g_{3D}\delta(\mathbf{r}-\mathbf{r^{\prime}})$ with strength $g_{3D}=4\pi\hbar^{2}a_{s}/m$ characterized by the $s$-wave scattering length, $a_{s}$, which can be positive ($a_{s}>0$) or negative ($a_{s}<0$) indicating that the effective interactions are repulsive or attractive, respectively. These scattering lengths can, depending on the atomic species, be tuned in magnitude and sign using magnetic [232, 233] or optical [234, 235] Feshbach resonances.